Determine the value of: 60 000 x 80 000 – 5 000 + 120 000 ÷ 15 000

4 800 000 000 – 5 000 + 8 = 4 799 995 008


1
Expert's answer
2021-05-18T14:23:20-0400

"60,000×80,000−5000+120,000÷15,000\\=4,800,000,000−5000+\frac{120,000}{15,000}\\=4,800,000,000−5000+8\\=4,800,000,000+8-5000\\=4,800,000,008-5000\\=4,799,995,008"